RC meeting and Run Plan, Saturday, December 10th
From SBS wiki
Return to the RC Meeting Calendar
Contents
(NO SATURDAY MEETING) RC Meeting Information
Current RC: | Eric King [ericking (@jlab.org or @temple.edu)] | |
Time: | Daily at 3:30PM (NO SATURDAY MEETING) | |
Location: | Counting House Conference Room |
Remote Connection Details
- Join ZoomGov Meeting
- https://jlab-org.zoomgov.com/j/1619032184?pwd=aWhTVWxya1ZZQzA1K2ZPSTd2ZX...
- Meeting ID: 161 903 2184
- Passcode: 260715
- Dial by your location
- +1 669 254 5252 US (San Jose)
- +1 646 828 7666 US (New York)
- +1 551 285 1373 US
- +1 669 216 1590 US (San Jose)
- 833 568 8864 US Toll-free
- Find your local number: https://jlab-org.zoomgov.com/u/acj3pTicMx
Shift Schedule
This RC Week
- Next empty shift is a Target Operator DAY shift on Sunday 12/11.
Upcoming RC Week
- There are now three (3) other remaining shifts to be filled before winter shutdown; I'm sure Andrew would like them filled asap so he doesn't have to chase down any shift workers.
Accelerator Status
- Accelerator is running smoothly.
- Hall C valve/vacuum repair is complete; they're coming back online Monday or Tuesday; they will be priority hall next week.
Current Status
We've been taking data on 3-He no major disruptions.
Beam Current
After much discussion the target group came to the decision yesterday afternoon/evening to increase the beam current from 36uA to 40uA utilizing the same size raster. Everything has been running smoothly since the change.
Target Polarization
Target polarization has been stable.
- 17:00 → 40.6% [4099678]
- 13:45 → 40.7% [4099629]
- 11:00 → 40.9% [4099588]
- 08:00 → 41.5% [4099536]
- 03:30 → 40.9% [4099455]
- 23:30 → 40.3% [4099385]
- 20:45 → 40.7% [4099346]
- 17:45 → 40.8% [4099308]
Issues Resolved
- The repeated DAQ issues which were difficult to resolve during Thursday evening into Friday morning seem to be resolved (at the moment).
- Bob ran into a minor issue in the early morning which cost about 15 minutes of data taking [4099495]
- The Moller HH magnet motor has remained functioning since Ellen did the restart on the power supply and is cooled down. Bill is looking into why the IOC for the Moller magnet is suffering a communication error with the alarm handler. [4099258]
- Hunter noticed the the laser software crashed [4099401]
Issues remaining
- BB GEM APV unreachable and cannot be repaired at this time.
Short Term Run Plan
- Continue production run on He3 target, 40 uA
- Use GEnII-noSBSGEMs DAQ config
- Use ps3=0, ps5=6, others =-1
- BBCal threshold at -366 mV, HCal threshold at -19 mV
- NMR measurement every 3 - 4 hours, close to 4 hours is better
Planned Accesses to the Hall
- I presume that repairs for the BB GEM APV will wait until shutdown.
Special Requests
Please email me ericking (@jlab.org or @temple.edu) if there's something to be added.
Notices in Effect for Shift Crew
- Power down SBS, BB, and Detector HV before making an access to the Hall
Data Collection Status
- 76 out of 141 days (~53.9%)
- (NEEDS TO BE UPDATED) Total uncorrected accumulated charge on He3: 51.782 C out of 232C (22.31%)
- 2-Pass: 13.467 C
- 3-Pass: 38.315 C (so far)